Fermi National Laboratory
 

Operation Summary for July 23, 2010 - July 30, 2010

 

7/23/10 – 7/30/10

LINAC

Shutdown Week 2

  • Preaccelerator
    • I- Source generator repair…………..50%
    • I- Source cleaning……………..……100%
    • I- Source heater supply upgrade…….50%
    • Both ion gauge replacements……...100%
  • Low Energy
    • LRF1 tuner/fingerstock repair………100%
    • LRF1 PA swap……………………….100%
    • Tanks 1 & 2 ion pump replacement...100%
    • Roughing pump maintenance……….100%
    • QPS card upgrade……………………50%
    • Modulator maintenance……….....…100%
    • Buncher RF coupler install……….....100%
    • LRF4 filament PLC upgrade………...100%
    • LRF5 PA filament leak repair……….100%
    • Modulator Acopian PS replacement…100%
  • High Energy
    • Solenoid flow meter upgrade………100%
    • KRF1 120V outlet install…………....100%
    • Distribution skid HX cleaning………..50%
    • Debuncher water skid pump relocation..100%
    • Debuncher water skid controls upgrade........50%
    • Transformer oil processing…………..60%
    • QPS maintenance…………………….25%
    • CS/PFN maintenance..........................…75%
    • SC tank 3 TLM install………………...60%
  • Electrical Work
    • LERF breaker panel replacement……..100%
    • LERF water skid wall disconnects….....100%
  • Controls Work
    • Pulse shifter chassis move…………...100%
    • LERF ACNET console installation…...100%
  • Next Week
    • We will use this week to finish all maintenance left to do on HE Linac
    • 7/30/10: Turn off all equipment to prepare for Linac power outage on Sunday
    • Monday: Start up H- Source
    • Tuesday: Turn on LE Linac after Master Sub outage in morning (thru LRF3)
    • Wednesday: Hopefully have beam available for NTF patient treatment
    • Friday: Begin turning on HE Linac.

Booster

Shutdown Week 2

  • Activities
    • Directional boring for new sump lines
      • Two new discharge lines have been installed
      • Booster pond connection has been made
    • BRF18 tuner replacement finishing up
    • BRF3 at MI60 awaiting testing
    • BRF15 tuner repair underway
    • Booster cable pulls finished
    • Drained tunnel LCW to replace ball valves
    • Next week install BRF1, BRF3, and BRF18

Antiproton Source

Rings/Transport

  • On Schedule

Beam Lines

  • On Schedule

Main Injector

MI Power Outages:

  • August 3rd 12:30 (30 minutes)
  • August 11th 7:00 (30 minutes)

Tunnel Work

MI-40

  • 12 new kicker cables pulled
  • Control/signal cables pulled
  • New cable tray installed
  • Flourinert supply and return lines installed in penetration
  • LCW supply and return lines installed in penetration
  • Leak checking kicker boxes
  • Leak checking beam pipe
  • Installed 7 of 7 loads in kickers

To-do:

  • Install LCW and flourinert in tunnel
  • Pressure test cooling lines
  • Remove Lead on the Lambertsons
  • Existing kicker cables put back
  • Cable tray replaced under cooling penetrations

MI-52 eCloud

  • New Carbon coated pipe installed
  • RFAs re-installed in new pipe
  • Pumping down

MI-8

  • 810 and 851 multiwires replaced
  • 813 ion pump replaced
  • Under vacuum/pumping down

MI-10

  • All three MI Ion Profile Monitors have new multi-channel plates
  • Back under vacuum

MI-30

  • MI-307 shielding stand delivered

MIRF:

  • Leak check – possible 606c vacuum problem
  • Small dripper to repair
  • Upstairs maintenance

Misc

  • Stainless beam pipe clamp replacement going quickly
  • Heat exchanger cleaning on schedule (7/11)
    • MI-60 Mag, RF equip, Cavity systems done,
    • MI-10, MI-20 & MI-62 are done. (30, 40, 50, 52 left)
  • MI-50 Upper and Lower transformer oil leaks repaired
  • Corona tested dry transformers
  • MI-10 two speed pond pump cable replacement scheduled for Monday

Recycler

Vertical Flying wire replacement (75%)

  • Still baking
    • Encountered some issues with the setup (+ power glitches)
    • Pressure was a bit high – Added one turbo on Wednesday

Replace beam valve solenoids (55%)

  • 9/16 done so far
    • Other areas will be done during the last week of the shutdown

Recycler vacuum maintenance

  • Leak check vacuum in 800 line – Next Monday
  • TSPs sublimation

Other

  • Power supplies maintenance
  • Controls nodes
  • Repair RF power amplifier on Station 3 – Testing should be taking place on Monday
  • Install shielding near BV308F – Next week

Electron Cooling

  • During closing today we found a problem when running the checklist, which might alter our plan to startup on Monday
  • Safety System Test on Tuesday morning
  • BNL personnel arrives on Wednesday for Pelletron studies

Inside Pelletron tank

  • Electrical work
    • Replaced SPD05 solenoid power supply
    • Replaced ion pump power supply
  • Mechanical work
    • Thorough cleaning of the chain (and the rest of the Pelletron)
    • LVDT (to monitor the chain elongation) installed yesterday
    • Replaced all conductions bands
    • Sheaves/chain/inductors aligned on Wednesday (NEC tech)

Outside Pelletron tank

  • Several ‘small’ maintenance tasks
    • SF6 Dryer/Recirculation system, SF6 transfer skid, Compressor…

MI-31 water system maintenance

  • Filters replaced and checked for debris
    • Strainers were clean
    • Only very small amount of debris when flushing

Remaining Tasks

  • Back flush SPC80 cooling solenoid
    • In the MI tunnel – At the end of the shutdown

Tevatron

Jobs for Next Week

  • Finish un rolling Tevatron magnets
  • Replace stands at D23-1 and E27-1
  • Lower D23-1 and E27-1 0.050 inches
  • Fix tilt meters B0 and D0 low beta quadrupoles
  • Reference E-0 pixel detector
  • Check B-1 for outside air leaks
  • Check D-2 for outside air leaks
  • Hipot of Main Ring magnets by TD

NuMI

Jobs for Next Week

  • Continue Surveys above the Target & Horn 1 in the Target Pile
  • Install New of Target Pile Fan and Motor
  • Install & Removal of Scaffolding for Target Pile Fan & Dehumidifier Repair
  • Re-Install Target Pile Fan & Dehumidification Duct Work
  • Box & Move Old Target Pile Fan & Motor to Site 40
  • Replace Plastic Air Lines on Target Pile Air Valves with Copper Line
  • Completed Replacement of Regeneration Motors on Dehumidifier Units A&B
  • Change Out Condensate Tanks Hoses
  • Clean Out the Stripline Cage Area
  • Replace Target Pile Humidity Monitors
  • Install Water Protection Over Absorber& Hadron Monitor
  • Repair PEI Units A & B at MI62
  • Assemble & Install TP Dehumidifier Service Platform
  • Completed Drain Cleaning in Underground Facility
    • Target Hall, Shaft Area, Support Rooms, Absorber Passageway, &Muon Alcoves
  • Upgrade Electrical Service & Replace Check Valve for Pre Target Sump Pumps
  • Investigate Increased Humidity Rates in Target Pile
    • Borascope East Floor Drain
  • Completed Replacement of Multiwires PW105, PW107, PW108, PW112
    • Completed Alignment Activities
    • Start Pump Down on Friday Afternoon
  • Continue Electrical Service Upgrade for ANU/NOvA in MI65 Electrical Room
  • Completed Cleaning of Target Pile Chiller (FESS)

Preview of Week #4

  • Continue with Target Pile Fan Repair
    • –Install Electrical Service
    • –Install Vibration Sensors & Instrumentation
    • –Balance Fan
  • Install Back draft Dampers on TP Dehumidifier Process Ducts
  • Munter Dehumidifier plant maintenance
  • Start installation for MINOS chiller & underground cooling loop upgrade
  • Start AH dehumidifier repair of heater controllers
  • Decay Pipe Concrete Tritium Samples for LBNE
  • Start Preparations for Target Change
    • –Remove & Clean R Blocks
    • –Remove Water Lines
    • –Complete Top of the Module As-Found Surveys in Chase

Cryogenic System

TeV Magnet Temperatures

status2

Cryo Operations and Maintenance Status

  • Tevatron at LHe temperatures except as noted above
  • 2 unplanned MSS power outages make stable cryogenic operations difficult. Helium inventory losses due to power outages replenished with gas jumbo high-pressure trailers; liquid helium to be delivered once full stable cryo operations is assured. CHL NRL plant off and we are relying on Linde deliveries (4-5 truckloads per day) to CHL’s 150,000 gallons of tank storage
  • Continued ringwide vacuum pump maintenance, repairs, and pumpdown of various components to assure stable Tevatron cryogenic system operations
  • 18 wet engine cold end overhauls are complete (8-10 hrs per unit), 1 overhaul in progress today
  • Magnet Kautzky relief flange clamps torques checked (several cold helium-atm leaks in past due to loose Marmon clamps); 10/24 houses done
  • FESS replaced the failed D0 water pump for the Mycom compressors
  • Ongoing numerous FRIG/Compr controls/instrumentation maintenance/repairs
  • Issues: F1 heat exchanger vacuum bad; E3 wet expander electrical problems

CHL HELIUM

inve

MiniBooNE

  • On Schedule

Collider Detector at Fermilab

  • On Schedule

D-Zero

  • On Schedule

Alignment

Week Two

Tevatron

  • Task: 5994: Level run around TeV ring all tie rods and floor plugs near D0 and B0 used for elevation
    • Work load:12 days
    • Date: Monday - Thursday July 26 2010,July 29 2010 – Completed Today
  • Task: 5998: Un-roll approximately 30 TeV quads and dipoles that have greater than 1 mR roll.
    • Work load:9 days
    • Date: Tuesday - Thursday July 26 2010,July 29 2010 – On Going

Switchyard

  • Task: 6025: 6013 needs a layout of Enclosure B for the referenced M-Septas currently located out at MI-65.
    • Work load: 1 days
    • Date: Monday July 26 2010 – Completed

MTA

  • Task: 6015: As-found the magnets upstream of UHB03 up to Q74
      • Realign UQ01
      • Re-align the permanent magnet if necessary
      • See Task Manager for details
    • Work load: .75 days
    • Date: Monday July 26 2010 – Completed

NuMI

  • Task: 5990: Alignment of 4 multiwires in NuMI MI section during the summer shutdown
    • Work load: 1.6 days
    • Start Date: Wednesday July 28 2010 – Completed

Week Three

Tevatron

  • Task: 5998: Un-roll approximately 35 TeV quads and dipoles that have greater than 1 mR roll.
    • Work load:9 days
    • Date: Monday - Thursday August 02 2010, August 06 2010– On Going
  • Task: 6019: Reference 2 new pixel detectors for Tevatron
    • Work load:1 days
    • Date: Wednesday August 04 2010
  • Task: 5964: Re-align quad D23-1 after stand replacement. Lower elevation by 0.050 inches from present value.
    • Work load: 0.5 days
    • Date: Monday - Thursday August 02 2010, August 06 2010

Booster

  • Task: 6033 Align three (3) RF cavities at Booster period L14 (BRF1) and Booster period L19 (BRF18) RF cavity at Booster period L15 (BRF3)..
    • Work load:1.5 days
    • Date: Tuesday - Wednesday August 03 2010, August 04 2010

Main Injector

  • Task: 6028 Level run of the Main Injector / Recycler tunnel
    • Work load:1.5 days
    • Date: Monday- Thursday August 02 2010, August 06 2010

Switchyard

  • Task:
    • Work load: 1.5 days
    • Date: Monday- Thursday August 02 2010, August 06 2010

Controls

Past Week

  • Recovery from unplanned Power Outages (Sat. & Wed.)
  • Applications and Device Database OS Upgrade
  • ACNET Console Restarts
  • Java Engine reboots
  • MI network fiber upgrade
  • TCLK $3E – SWYD beam permit drop
  • General Controls Maintenance

Upcoming Week

  • Saturday & Sunday - ACNET Down
  • CUB work & Cross Gallery power outage
  • Network switch maintenance
  • General Controls Maintenance & Support

Switchyard 120

  • EE Support connected the high voltage cable to the Meson Separators
  • The supply was verified by interlock personnel that it is tied to the Enclosure B Safety System
  • An alignment has been completed by surveyors on both Meson Separators
  • Controls have reused the existing cables for motor controls and LVDT readbacks
  • After the power outage on Wednesday for Feeder 43, the Enclosure C Upstream sump pump failed
  • FESS personnel investigated and found two problems
    • The discharge for one pump was broken
    • The other pump’s float was tangled between the two pumps not allowing the pump to trigger
  • Reminder
    • Enclosure B doors will be changed back from the Tevatron shutdown key, to the original Enclosure B on Friday, 7/30/10

FESS

Electrical

  • Feeder 46A Moved to August 5th
  • Master Substation Back – ON “Monday”
  • Kautz Road Substation – OFF “Tuesday”
  • Feeder 24 Pbar OFF 4 to 6 hours “Wednesday”
  • ODH Fan Repairs
  • Feeder 44 Partial Outage SCI-ED, Mini Boone, MINOS “Today”

Mechanical

  • Sump Pump Replacement @ SY ENC “C”
  • Sump Pump Replacement @ AO10, AP0TH
  • Exhaust Fan Repair @ AP0TH
  • TEV Pond Pumps “repack”

HVAC

  • MI – Repair DesertAire
  • BOOSTER – Economizer on B288
  • MI – Revisit MarvAires in construction areas
  • Linac Pre Acc EDPac unit – Humidistat needs replaced “parts

Operations Engineering

  • CUB - Pipe Inspection
  • Booster Sump Pump Upgrade
  • Dual Temp Piping Modification “near completion”
  • Switch Yard Service Building TEE
  • TEV – High Silt level extensions A0, C0, D0, D3, F0, RF
  • X-Gallery Piping Work “this weekend”

CUB

  • LCW off to Pbar & Booster heat exchangers @ 1 PM today to replace leaking tower valves
  • CUB Shutdown all set for Saturday & Sunday
  • Chiller Epoxy coatings near completion
  • Chiller #5 control transformer Repair (4160 V outage)

Findings

  • High Silt levels at the suction boxes
  • Zebra Mussels found in simplex basket strainers

Safety

Remember to utilize Take 5 principles to Reduce Injuries and Environmental Impacts

Radiation Safety

  • Beam Permits and Run Conditions have been rescinded for all machines except Linac, NTF, & Pelletron Diagnostic Mode
    • Accelerator restart sheets are available
  • Enclosure B will be re-cored to the enclosure access keys after 3:00 pm today to allow for separator conditioning
    • Keys for access to Enclosure B will be issued daily from the MCR
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/02/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory